OpenStudy (anonymous):

On a scale drawing showing the floor plan of an apartment, the rectangular living room measures 3.5 cm by 6 cm. If 1 centimeter represents 1.5 meters, what is the area of the floor of the actual living room, in square meters? Express your answer as a decimal to the nearest hundredth.

OpenStudy (anonymous):

3.5 * 1.5 = 5.25 m 6 * 1.5 = 9 m 5.25 * 9 = 47.25 m^2

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

convert 3.5 and 6 to the actual living room by multiplying 6. Then you multiply both of them to get the area

OpenStudy (anonymous):

Ok